# EPOCHS IV: SED Modelling Assumptions and their impact on the Stellar Mass Function at using PEARLS and public JWST observations

### v1

@@ Abstract
We utilize deep JWST NIRCam observations for the first direct constraints on the Galaxy Stellar Mass Function (GSMF) at . Our EPOCHS v1 sample includes 1120 galaxy candidates at taken from a consistent reduction and analysis of publicly available deep JWST NIRCam data covering the PEARLS, CEERS, GLASS, JADES GOOD-S, NGDEEP, and SMACS0723 surveys, totalling 187 arcmin. We investigate the impact of SED fitting methods, assumed star formation histories (SFH), dust laws, and priors on galaxy masses and the resultant GSMF. Whilst our fiducial GSMF agrees with the literature at , we find that the assumed SFH model has a large impact on the GSMF and stellar mass density (SMD), finding a 0.75 dex increase in the SMD at between a flexible non-parametric and standard parametric SFH. Overall, we find a flatter SMD evolution at than some studies predict, suggesting a rapid buildup of stellar mass in the early Universe. We find no incompatibility between our results and those of standard cosmological models, as suggested previously, although the most massive galaxies may require a high star formation efficiency. We find that the ‘Little Red Dot’ galaxies dominate the GSMF at high-masses, necessitating a better understanding of the relative contributions of AGN and stellar emission. We show that assuming a theoretically motivated top-heavy IMF reduces stellar mass by 0.5 dex without affecting fit quality, but our results remain consistent with existing cosmological models with a standard IMF.
@@ 1 Introduction
An immediate result was a potential overabundance of high-redshift galaxies compared to theoretical predictions and extrapolations of Hubble Space Telescope (HST)/Spitzer results (Naidu et al., 2022; Haslbauer et al., 2022; Mauerhofer & Dayal, 2023; Mason et al., 2023). The mere existence of some of the galaxies at the inferred redshifts and stellar masses, ( M at ) seem to be in tension with standard CDM cosmology given the small areas currently probed with JWST (Labbé et al., 2023; Lovell et al., 2023; Boylan-Kolchin, 2023). However, spectroscopic observations of some of the highest-mass candidates in Labbé et al. (2023) have reduced their redshifts and stellar masses, or hinted at the presence of an Active Galactic Nuclei (AGN), and hence reduced possible CDM tension (Kocevski et al., 2023; Fujimoto et al., 2022).
Initial overlap of galaxy candidates between independent studies was poor, but has since improved due to agreement on photometric calibration and reduction techniques (Adams et al., 2023; Rieke et al., 2022). Spectroscopic confirmation exists only for a fraction of potential candidates, but most spectroscopic programs have had a high success rate, along with a few notable low- interlopers (Robertson et al., 2023b; Curtis-Lake et al., 2023; Arrabal Haro et al., 2023; Tang et al., 2023; Laseter et al., 2023; Fujimoto et al., 2023; Bunker et al., 2023a; Wang et al., 2023a).
@@
Another assumption is that of a possible universal initial mass function (IMF), which predicts the number of stars as a function of stellar mass (Salpeter, 1955). The presence of low-mass stars, which dominate the stellar mass, is inferred entirely from the shape of the assumed IMF in most galaxies. A universal IMF (e.g. Salpeter, 1955; Chabrier et al., 2000; Kroupa, 2001), which is constant across time and space, has long been assumed in the majority of galaxy studies. Models of the the physics of high- star formation suggest that it may have deviated from the universal IMF above redshift (Hopkins et al., 2005; Jermyn et al., 2018; Steinhardt et al., 2021, 2023). Low metallicity, high star formation rates, an increasing CMB temperature and high cosmic ray density could all contribute to heating of star formation regions and lead to a top-heavy IMF at high redshift (Gunawardhana et al., 2011; Clauwens et al., 2016; Papadopoulos et al., 2011). Observations of local elliptical galaxies have also found them to be inconsistent with the universal IMF, and instead find evidence for an IMF gradient, with evidence for a different IMF between younger and older stellar populations (e.g. Weidner et al., 2013, & references therein). Studies such as Steinhardt et al. (2023); Sneppen et al. (2022); Woodrum et al. (2023) have introduced temperature-dependent IMF models for use in SED-fitting, which can decrease stellar mass estimates by up to 1 dex at constant redshift. Constraining the IMF is extremely difficult, but recent studies are beginning to find possible evidence for a top-heavy IMF in the early Universe (e.g. Katz et al., 2022; Cameron et al., 2023; Mowla et al., 2024). Ultimate conclusions on this are however far from certain.
In the EPOCHS paper series we have presented an independent and consistent reduction of deep JWST observations from available GTO, GO and ERS data including the CEERS, GLASS, SMACS-0723, JADES and PEARLS fields. We found 1165 robust galaxy candidates above redshift , with a total area of 187 arcmin (Conselice et al., in prep; Adams et al., 2023).
@@ 2 Products and Data Reduction
This section briefly details the JWST programs and data products used in this analysis. For further details on the fields used please see Austin et al. (2023), Adams et al. (2023) and Conselice et al. (in prep). Table 1 shows the available unmasked areas, JWST NIRCam filters and 5 depths of each dataset used.
@@ 2.1 PEARLS
We incorporate NIRCam observations from the proprietary GTO survey Prime Extragalactic Areas for Reionization Science (PEARLS, PI: R. Windhorst & H.Hammel, PID: 1176 & 2738, Windhorst et al., 2023). We use observations of three fields targeting gravitationally lensed clusters, and one blank field consisting of a mosaic of 8 JWST NIRCam pointings. The gravitationally lensed clusters consist of MACS J0416.1-2403 (hereafter referred to as MACS-0416), El Gordo (Menanteau et al., 2012, , ACT-CL J0102-4915 in the Atacama Cosmology survey, [) and Clio (, Designation GAMA100050 in the GAMA Galaxy Group Catalog v6+, Robotham et al., 2011). El Gordo and Clio have been visited once with JWST/NIRCam, with the cluster centered in one NIRCam module and the other observing a neighbouring blank field away (Griffiths et al., 2018). MACS-0416 has been observed 3 times, resulting in 3 separate parallel observations at different position angles. The PEARLS blank field is the North Ecliptic Pole Time Domain Field (Jansen & Windhorst, 2018, (NEP-TDF,). The NEP-TDF is positioned so it can be observed throughout the year, making it ideal for time-domain science and constructing a large deep field. Observations of the NEP-TDF consist of four pairs of overlapping NIRCam pointings (8 pointings total), with each of these four pairs orientated at 90 degree intervals like spokes on a windmill. NIRCam observations of the NEP-TDF, El Gordo and MACS0416 use the standard 8 photometric bands; F090W, F115W, F150W, F200W, F277W, F356W, F410M and F444W. Clio uses 6 of the 8 previous bands, but lacks F115W and F410M. Within the NEP-TDF field we incorporate HST Advanced Camera for Surveys Wide Field Channel (ACS/WFC) imaging in the F606W filter, collected as part of the GO-15278 (PI: R. Jansen) and GO-16252/16793 (PIs: R. Jansen & N. Grogin) between October 1 2017 and October 31 2022. Mosaics of these data, astrometrically aligned to Gaia/DR3 and resampled on 003 pixels, were made available pre-publication by R. Jansen & R. O’Brien (private comm. O’Brien et al., 2024). For full details of the PEARLS program please see Windhorst et al. (2023).
@@ 3 Catalogue Creation and Sample Selection
Full details of our catalogue creation and sample selection pipeline, called GALFIND, is available in Conselice et. al. (in prep). We briefly summarise the procedure used here.
@@ 3.2 Sample Selection
Our selection criteria are similar to other high- galaxy studies, such as Hainline et al. (2023); Finkelstein et al. (2023a, b); Naidu et al. (2022); Castellano et al. (2022), who also fit galaxies using EAZY-py, and select their samples from the rest-UV SNR of their candidates and the resultant redshift PDF and from the SED fitting. SNR requirements vary between studies, and are also somewhat dependent on the size of the extraction aperture used. Our 016 radius apertures are larger than the 01 radius apertures used in Finkelstein et al. (2023b), resulting in a number of their candidates being removed from our sample because they do not meet our 5 threshold, despite agreeing on the photo- solution. Other studies, such as as Harikane et al. (2023); Yan et al. (2022), use a combination of color-color cuts and SED-fitting. A comprehensive comparison of the EPOCHS v1 galaxy sample to other studies is given in Appendix A of Adams et al. (2023).
@@ 4 Galaxy Properties from SED Fitting
The basic properties, redshift distribution, number counts and UVJ colors of the EPOCHS v1 sample are described in detail in Conselice et al. (in prep). The UV properties (, slopes) of this sample are explored in Austin et al. (in prep). The UV luminosity function is presented in Adams et al. (2023), Here we briefly summarise the basic statistics of the EPOCHS v1 sample, following the selection criteria described in § 3.2.
@@
Across the fields described in § 2, 1214 galaxies pass our initial selection criteria with over a total unmasked area of 187.27 arcmin, comprising one of the largest samples of high- JWST-selected galaxies. Our visual inspection removes 49 completely, leaving 1165 galaxies, which we refer to as the EPOCHS v1 sample. Of these 1165, 1054 are classed as “certain” and 111 “uncertain” by our visual inspection. In this work we choose to include the visually “uncertain” candidates, in order to avoid potentially underestimating the stellar mass function. Filtering the sample to the redshift range used in this work, our fiducial sample consists of 1120 galaxies.
@@
The next sections detail the Bagpipes and Prospector SED fitting performed for all galaxies in the EPOCHS v1 sample. We perform this SED fitting only for our EAZY-py-selected sample. The purpose of this SED-fitting is to analyse the properties of the stellar and nebular components of these high- galaxies. We note that whilst the majority of galaxies in the EPOCHS v1 sample do not appear significantly extended beyond our SExtractor extraction aperture, all stellar masses quoted in the following results have been corrected by the ratio between the SExtractor FLUX_AUTO_F444W and the aperture corrected FLUX_APER_F444W fluxes, if it exceeded unity, to account for any residual flux outside the aperture. We use the longest wavelength band to correct stellar masses as this is most representative of the rest-frame optical emission.
@@ 4.1 Bagpipes
Bagpipes (Bayesian Analysis of Galaxies for Physical Inference and Parameter EStimation, Carnall et al., 2018, 2019)) is a Python package which uses Bayesian methods to fit galaxy SEDs to photometry. Our fiducial Bagpipes run uses the Bruzual & Charlot (2003) 2016 stellar population synthesis models, and a Kroupa (2001) IMF.
@@
Bagpipes can construct and fit SEDs with a variety of SFH models, dust models and priors. We perform multiple fits for each galaxy in order to test the consistencies in derived galaxy parameters. For simplicity our approach is to define an initial ‘fiducial’ model, and then swap out individual model components or priors for other choices. Our model parameters, priors and hyper-parameters are detailed in Table 2. This is similar to the approach of Furtak et al. (2021), who define a reference model which is then compared to alternative models. In the rest of this section we further explain our choices of models and priors.
@@
For the star formation history we test both parametric and non-parametric models, which have been shown to impact the stellar mass estimates (Leja et al., 2019; Tacchella et al., 2022). For our fiducial model, we use a parametric lognormal SFH which allows us to recreate the rising SFHs we expect in the early Universe. We compare this SFH with another commonly used parametrization, the delayed exponential. Details of the implementation of these parametric SFHs are available in Carnall et al. (2019). We also test a non-parametric “continuity” SFH model similar to the model added to Prospector in Leja et al. (2019). This SFH model fits the star formation rates in fixed time bins, with SFR between bins linked by a Student’s t-distribution. We recreate the methodology of Tacchella et al. (2022), by fitting both a “continuity” model, where the Student’s t-distribution has hyper-parameters = 0.3 and = 2, which weights against rapid changes in star formation rate, as well as a “continuity bursty” model, with = 1 and = 2, which allows more stochastic star formation, with rapid bursts and quenching more similar to the SFH inferred at high-. As is done in Tacchella et al. (2022) we fit 6 SFH bins for both models, with the first bin fixed to a lookback time of 0-10 Myr, the last bin ending at z = 20, and the other 4 bins are equally log-spaced in lookback time. Leja et al. (2019) showed that this model is relatively insensitive to the number of bins used, as long as the number exceeds 4. Like Tacchella et al. (2022), we assume there is no star formation at (Tacchella et al., 2022; Bowman et al., 2018; Jaacks et al., 2019). It is important to note that this model allows but does not require ‘bursty’ SFHs, and smooth or quenched SFHs are also possible if favoured during the fitting.
@@
We include emission lines and nebular continuum based on CLOUDY v17.03 (Ferland et al., 2017). We regenerate CLOUDY models in order to probe a wider range of the ionisation parameter U between -3 -1 using a CLOUDY configuration file distributed with Bagpipes. For the ionization parameter we use a uniform prior (in log space).
@@
| Common Parameters | ||
| Parameter | Prior/Value (Min, Max) | Description |
| EAZY-py Posterior PDF () | Redshift | |
| SPS Model | Bruzual & Charlot (2003); BPASS v2.2.1 | Stellar population synthesis model |
| IMF | Kroupa (2001); default BPASS IMF | Stellar Initial Mass Function |
| Dust Law parametrization | Calzetti et al. (2000); Salim et al. (2018) | Dust law |
| Charlot & Fall (2000) | ||
| A | Log-uniform: (, 10); uniform (0, 6) | V-band attenuation (all stars) |
| SFH | lognormal; “continuity bursty”; delayed- | Star formation history |
| uniform: (5, 12) | Surviving stellar mass | |
| Z | log-uniform: (0.005, 5); uniform (0, 3) | stellar metallicity |
| Z | Fixed to Z | gas-phase metallicity |
| uniform: (-3, -1) | Ionization Parameter |
| Model Specific Parameters | |||
| Model | Parameter | Prior/Value (Min, Max) | Description |
| Fiducial | t | uniform: (10 Myr, 15 Gyr) | Age of Universe at peak SFR |
| FWHM | uniform: (10 Myr, 15 Gyr) | FWHM of SFH | |
| delayed- SFH | uniform: (10 Myr, 15 Gyr) | e-folding timescale | |
| Age | log-uniform: (10 Myr, t | Time since SF began | |
| “‘continuity bursty”’ | N | 6 bins (5 fitted parameters) | First bin 0 – 10 Myr, SF begins at z = 20, |
| non-parametric SFH | others distributed equally in log lookback time | ||
| d | Student’s-t: | Ratio of SFR in adjacent bins, coupled by | |
| Charlot & Fall (2000) | n | clipped normal: | Power-law slope of attenuation curve (A ) |
| Dust Law | (0.3, 2.5) | For Calzetti et al. (2000) n | |
| clipped normal: | A/A ratio between young and old stars | ||
| (1, 3) | |||
| Salim et al. (2018) Dust Law | clipped normal: | Deviation from Calzetti et al. (2000) slope | |
| (-0.3, 0.3) | |||
| uniform (0, 5) | Strength of 2175Å bump | ||
| BPASS SPS Model | No additional components | ||
| Uniform A Prior | No additional components | ||
| Uniform Z Prior | No additional components |

### v3

@@ Abstract
We utilize deep JWST NIRCam observations for the first direct constraints on the Galaxy Stellar Mass Function (GSMF) at . Our EPOCHS v1 sample includes 1120 galaxy candidates at taken from a consistent reduction and analysis of publicly available deep JWST NIRCam data covering the PEARLS, CEERS, GLASS, JADES GOOD-S, NGDEEP, and SMACS0723 surveys, totalling 187 arcmin2. We investigate the impact of SED fitting methods, assumed star formation histories (SFH), dust laws, and priors on galaxy masses and the resultant GSMF. Whilst our fiducial GSMF agrees with the literature at , we find that the assumed SFH model has a large impact on the GSMF and stellar mass density (SMD), finding a 0.75 dex increase in the SMD at between a flexible non-parametric and standard parametric SFH. Overall, we find a flatter SMD evolution at than some studies predict, suggesting a rapid buildup of stellar mass in the early Universe. We find no incompatibility between our results and those of standard cosmological models, as suggested previously, although the most massive galaxies may require a high star formation efficiency. We find that the ‘Little Red Dot’ galaxies dominate the GSMF at high-masses, necessitating a better understanding of the relative contributions of AGN and stellar emission. We show that assuming a theoretically motivated top-heavy IMF reduces stellar mass by 0.5 dex without affecting fit quality, but our results remain consistent with existing cosmological models with a standard IMF.
@@ 1 Introduction
An immediate result was a potential overabundance of high-redshift galaxies compared to theoretical predictions and extrapolations of Hubble Space Telescope (HST)/Spitzer results (Naidu et al., 2022; Haslbauer et al., 2022; Mauerhofer & Dayal, 2023; Mason et al., 2023). The mere existence of some of the galaxies at the inferred redshifts and stellar masses, ( M⊙ at ) seem to be in tension with standard CDM cosmology given the small areas currently probed with JWST (Labbé et al., 2023; Lovell et al., 2023; Boylan-Kolchin, 2023; Desprez et al., 2024). However, spectroscopic observations of some of the highest-mass candidates in Labbé et al. (2023) have reduced their redshifts and stellar masses, or hinted at the presence of an Active Galactic Nuclei (AGN), and hence reduced possible CDM tension (Kocevski et al., 2023; Fujimoto et al., 2022). Massive () quiescent galaxies at also appear to challenge theories of galaxy evolution given the old stellar populations implied by their star formation histories (Glazebrook et al., 2024; Casey et al., 2024; Carnall et al., 2024).
Initial overlap of galaxy candidates between independent studies was poor, but has since improved due to agreement on photometric calibration and reduction techniques (Rieke et al., 2022; Adams et al., 2023). Spectroscopic confirmation exists only for a fraction of potential candidates, but most spectroscopic programs have had a high success rate, along with a few notable low- interlopers (Robertson et al., 2023b; Curtis-Lake et al., 2023; Arrabal Haro et al., 2023; Tang et al., 2023; Laseter et al., 2023; Fujimoto et al., 2023; Bunker et al., 2023a; Wang et al., 2023a).
@@
Another assumption is that of a possible universal initial mass function (IMF), which predicts the number of stars as a function of stellar mass (Salpeter, 1955). The presence of low-mass stars, which dominate the stellar mass, is inferred entirely from the shape of the assumed IMF in most galaxies. A universal IMF (e.g. Salpeter, 1955; Chabrier et al., 2000; Kroupa, 2001), which is constant across time and space, has long been assumed in the majority of galaxy studies. Models of the the physics of high- star formation suggest that it may have deviated from the universal IMF above redshift (Hopkins et al., 2005; Jermyn et al., 2018; Steinhardt et al., 2021, 2023). Low metallicity, high star formation rates, an increasing CMB temperature and high cosmic ray density could all contribute to heating of star formation regions and lead to a top-heavy IMF at high redshift (Gunawardhana et al., 2011; Clauwens et al., 2016; Papadopoulos et al., 2011; Cueto et al., 2024). Observations of local elliptical galaxies have also found them to be inconsistent with the universal IMF, and instead find evidence for an IMF gradient, with evidence for a different IMF between younger and older stellar populations (e.g. Weidner et al., 2013; Yan et al., 2021, & references therein). Studies such as Jeřábková et al. (2018); Steinhardt et al. (2023); Sneppen et al. (2022); Woodrum et al. (2023); Cueto et al. (2024) have introduced temperature/time-dependent IMF models for use in SED-fitting, which can decrease stellar mass estimates by up to 1 dex at constant redshift. Constraining the IMF is extremely difficult, but recent studies are beginning to find possible evidence for a top-heavy IMF in the early Universe (e.g. Katz et al., 2022; Cameron et al., 2023; Mowla et al., 2024). Ultimate conclusions on this are however far from certain.
In the EPOCHS paper series we have presented an independent and consistent reduction of deep JWST observations from available GTO, GO and ERS data including the CEERS, GLASS, SMACS-0723, JADES and PEARLS fields. We found 1165 robust galaxy candidates above redshift , with a total area of 187 arcmin2 (Conselice et al., 2024; Adams et al., 2023).
@@ 2 Products and Data Reduction
This section briefly details the JWST programs and data products used in this analysis. For further details on the fields used please see Austin et al. (2023), Adams et al. (2023) and Conselice et al. (2024). Table 1 shows the available unmasked areas, JWST NIRCam filters and 5 depths of each dataset used.
@@ 2.1 PEARLS
We incorporate NIRCam observations from the proprietary GTO survey Prime Extragalactic Areas for Reionization Science (PEARLS, PI: R. Windhorst & H.Hammel, PID: 1176 & 2738, Windhorst et al., 2023). We use observations of three fields targeting gravitationally lensed clusters, and one blank field consisting of a mosaic of 8 JWST NIRCam pointings. The gravitationally lensed clusters consist of MACS J0416.1-2403 (hereafter referred to as MACS-0416), El Gordo (, ACT-CL J0102-4915 in the Atacama Cosmology survey, (Menanteau et al., 2012) and Clio , Designation GAMA100050 in the GAMA Galaxy Group Catalog v6+, Robotham et al., 2011). El Gordo and Clio have been visited once with JWST/NIRCam, with the cluster centered in one NIRCam module and the other observing a neighbouring blank field away (Griffiths et al., 2018). MACS-0416 has been observed 3 times, resulting in 3 separate parallel observations at different position angles. The PEARLS blank field is the North Ecliptic Pole Time Domain Field (Jansen & Windhorst, 2018, (NEP-TDF,). The NEP-TDF is positioned so it can be observed throughout the year, making it ideal for time-domain science and constructing a large deep field. Observations of the NEP-TDF consist of four pairs of overlapping NIRCam pointings (8 pointings total), with each of these four pairs orientated at 90 degree intervals like spokes on a windmill. NIRCam observations of the NEP-TDF, El Gordo and MACS0416 use the standard 8 photometric bands; F090W, F115W, F150W, F200W, F277W, F356W, F410M and F444W. Clio uses 6 of the 8 previous bands, but lacks F115W and F410M. Within the NEP-TDF field we incorporate HST Advanced Camera for Surveys Wide Field Channel (ACS/WFC) imaging in the F606W filter, collected as part of the GO-15278 (PI: R. Jansen) and GO-16252/16793 (PIs: R. Jansen & N. Grogin) between October 1 2017 and October 31 2022. Mosaics of these data, astrometrically aligned to Gaia/DR3 and resampled on 003 pixels, were made available pre-publication by R. Jansen & R. O’Brien (private comm. O’Brien et al., 2024). For full details of the PEARLS program please see Windhorst et al. (2023).
@@ 3 Catalogue Creation and Sample Selection
Full details of our catalogue creation and sample selection pipeline, called GALFIND, is available in Conselice et al. (2024). We briefly summarise the procedure used here.
@@ 3.1 Catalogue Creation
We choose not to PSF homogenize our imaging, and instead rely on a band dependent aperture correction to account for the changing PSF. Other studies have shown that the WebbPSF can underestimate the broadening of the PSF at small scales (e.g. Morishita et al., 2024; Weaver et al., 2024), and measuring an empirical PSF is extremely difficult in some of our fields given both the limited number of isolated stars in single pointing fields (e.g. NGDEEP, El Gordo or CLIO), or more complex position dependent PSFs due to the stacking of exposures taken at different times and position angles (e.g. NEP and MACS-0416). In the JADES-Deep-GS field we tested whether not PSF-matching had a systematic impact on the derived fluxes, or the galaxy properties derived from SED fitting. We repeated our full analysis of this field, using PSF matched imaging, using PSF models and kernels derived following the methodology of Weaver et al. (2024); Suess et al. (2024), finding very little impact ( 0.05 magnitudes) on the resultant aperture fluxes across all bands, which is well within the 10% flux uncertainty floor used when SED fitting.
@@ 3.2 Sample Selection
Our selection criteria are similar to other high- galaxy studies, such as Hainline et al. (2023); Finkelstein et al. (2023a, b); Naidu et al. (2022); Castellano et al. (2022), who also fit galaxies using EAZY-py, and select their samples from the rest-UV SNR of their candidates and the resultant redshift PDF and from the SED fitting. The requirement was chosen to ensure that the high-z solution is a better fit than any possible low-z explanation, and is consistent with the criteria used in other studies, such as Finkelstein et al. (2023b). SNR requirements vary between studies, and are also somewhat dependent on the size of the extraction aperture used. Our 016 radius apertures are larger than the 01 radius apertures used in Finkelstein et al. (2023b), resulting in a number of their candidates being removed from our sample because they do not meet our 5 threshold, despite agreeing on the photo- solution. Other studies, such as as Harikane et al. (2023); Yan et al. (2022), use a combination of color-color cuts and SED-fitting. A comprehensive comparison of the EPOCHS v1 galaxy sample to other studies is given in Appendix A of Adams et al. (2023).
@@ 4 Galaxy Properties from SED Fitting
The basic properties, redshift distribution, number counts and UVJ colors of the EPOCHS v1 sample are described in detail in Conselice et al. (2024). The UV properties (, slopes) of this sample are explored in Austin et al. (2024). The UV luminosity function is presented in Adams et al. (2023), Here we briefly summarise the basic statistics of the EPOCHS v1 sample, following the selection criteria described in § 3.2.
For full transparency, we list the number of candidate objects removed by each stage of our selection criteria as follows. We note that as we apply all our selection criteria to all galaxies, each object can fail multiple selection criteria. Across all the survey fields used, the parent sample consists of 211,469 objects which are unmasked in all available photometric bands. Of these, only 1545 have photometric redshifts between and also pass the SNR criteria (criteria 1 - 3, above). 32 are removed because they fail criteria 4, as the photo-z is not well constrained. 9 are removed by criteria 5, as the suggests that the best fitting SED does not reproduce the observed photometry. 333 are removed by criteria 6, which requires that the high- solution has a robust statistical improvement when compared with the low- () solution. 63 brown dwarfs are identified and removed from the sample in criteria 7, which is discussed further in § 5.2.2. Finally 5 hot-pixels are removed by criteria 8, which have unphysically small sizes. Accounting for the 49 galaxies which fail multiple of the preceding criteria, this leaves 1214 galaxies which pass our initial selection criteria with over a total unmasked area of 187.27 arcmin2, comprising one of the largest samples of high- JWST-selected galaxies.
Our requirement constraining the redshift PDF is important for confident photo-’s and does not significantly bias our sample as only 7 galaxies of the full parent sample are rejected for this reason alone, and visual inspection of these 7 objects show they have SNR 10 and physically implausible SEDs.
@@
Our visual inspection removes 49 completely, leaving 1165 galaxies, which we refer to as the EPOCHS v1 sample. Of these 1165, 1054 are classed as “certain” and 111 “uncertain” by our visual inspection. In this work we choose to include the visually “uncertain” candidates, in order to avoid potentially underestimating the stellar mass function. Filtering the sample to the redshift range used in this work, our fiducial sample consists of 1120 galaxies.
@@
The next sections detail the Bagpipes and Prospector SED fitting performed for all galaxies in the EPOCHS v1 sample. We perform this SED fitting only for our EAZY-py-selected sample. The purpose of this SED-fitting is to analyse the properties of the stellar and nebular components of these high- galaxies. We note that whilst the majority of galaxies in the EPOCHS v1 sample do not appear significantly extended beyond our SExtractor extraction aperture, all stellar masses quoted in the following results have been corrected by the ratio between the SExtractor FLUX_AUTO_F444W and the aperture corrected FLUX_APER_F444W fluxes, if it exceeded unity, to account for any residual flux outside the aperture. We use the longest wavelength band to correct stellar masses as this is most representative of the rest-frame optical emission. We choose to perform all SED fitting using aperture fluxes, which ensures flux is measured from the same aperture in all filters, and avoids catastrophic issues with the SExtractor Kron aperture estimates, which sometimes become unphysically small or large. Some studies (e.g. Suess et al., 2024; Weaver et al., 2024; Weibel et al., 2024) calculate total fluxes by scaling all the overall normalisation of all fluxes by the ratio between a Kron and aperture flux in a detection or reference band, which has no impact on the inferred galaxy colors. As SED fitting depends on the measured colors to infer the star formation history and other parameters, and the total stellar mass simply scales linearly with the normalization. Our tests using the JADES-Deep-GS field, in which we repeat our full catalogue creation, selection and SED fitting procedure on PSF-matched imaging with ’TOTAL’ fluxes, finds resultant stellar masses consistent to within a mean (median) offset of 0.03 (0.01) dex. Given that this is well within the statistical uncertainty of our stellar mass estimates, we find that our methodology of using a per band aperture correction and aperture corrected stellar masses is completely consistent with the results of comparable studies.
@@ 4.1 Bagpipes
Bagpipes (Bayesian Analysis of Galaxies for Physical Inference and Parameter EStimation, Carnall et al., 2018, 2019) is a Python package which uses Bayesian methods to fit galaxy SEDs to photometry. Our fiducial Bagpipes run uses the Bruzual & Charlot (2003) 2016 stellar population synthesis models, and a Kroupa (2001) IMF.
@@
Bagpipes can construct and fit SEDs with a variety of SFH models, dust models and priors. We perform multiple fits for each galaxy in order to test the consistencies in derived galaxy parameters. For simplicity our approach is to define an initial ‘fiducial’ model, and then swap out individual model components or priors for other choices. Our model parameters, priors and hyper-parameters are detailed in § 4.1. This is similar to the approach of Furtak et al. (2021), who define a reference model which is then compared to alternative models. In the rest of this section we further explain our choices of models and priors.
@@
For the star formation history we test both parametric and non-parametric models, which have been shown to impact the stellar mass estimates (Leja et al., 2019; Tacchella et al., 2022). For our fiducial model, we use a parametric lognormal SFH which allows us to recreate the rising SFHs we expect in the early Universe. We compare this SFH with another commonly used parametrization, the delayed exponential. Details of the implementation of these parametric SFHs are available in Carnall et al. (2019). We also test a non-parametric “continuity” SFH model similar to the model added to Prospector in Leja et al. (2019). This SFH model fits the star formation rates in fixed time bins, with SFR between bins linked by a Student’s t-distribution. We recreate the methodology of Tacchella et al. (2022), by fitting both a “continuity” model, where the Student’s t-distribution has hyper-parameters and , which weights against rapid changes in star formation rate, as well as a “continuity bursty” model, with and , which allows more stochastic star formation, with rapid bursts and quenching more similar to the SFH inferred at high-. As is done in Tacchella et al. (2022) we fit 6 SFH bins for both models, with the first bin fixed to a lookback time of 0-10 Myr, the last bin ending at z = 20, and the other 4 bins are equally log-spaced in lookback time. Leja et al. (2019) showed that this model is relatively insensitive to the number of bins used, as long as it is more than four. Like Tacchella et al. (2022), we assume there is no star formation at (Tacchella et al., 2022; Bowman et al., 2018; Jaacks et al., 2019). It is important to note that this model allows but does not require ‘bursty’ SFHs, and smooth or quenched SFHs are also possible if favoured during the fitting.
@@
We include emission lines and nebular continuum based on CLOUDY v17.03 (Ferland et al., 2017). We regenerate CLOUDY models in order to probe a wider range of the ionisation parameter U between -3 -1 using a CLOUDY configuration file distributed with Bagpipes.
@@
cell11 = c=2,
Common Parameters &
Parameter Prior/Value (Min, Max) Description
EAZY-py Posterior PDF () Redshift
SPS Model Bruzual & Charlot (2003); BPASS v2.2.1 Stellar population synthesis model
IMF Kroupa (2001); default BPASS IMF Stellar Initial Mass Function
Dust Law parametrization Calzetti et al. (2000); Salim et al. (2018) Dust law
Charlot & Fall (2000)
A Log-uniform: (, 10); uniform (0, 6) V-band attenuation (all stars)
SFH lognormal; “continuity bursty”; delayed- Star formation history
uniform: (5, 12) Surviving stellar mass
Z log-uniform: (0.005, 5); uniform (0, 3) stellar metallicity
Z Fixed to Z⋆ gas-phase metallicity
uniform: (-3, -1) Ionization Parameter
cell11 = c=2
Model Specific Parameters
Model Parameter Prior/Value (Min, Max) Description
Fiducial t uniform: (10 Myr, 15 Gyr) Age of Universe at peak SFR
FWHM uniform: (10 Myr, 15 Gyr) FWHM of SFH
delayed- SFH uniform: (10 Myr, 15 Gyr) e-folding timescale
Age log-uniform: (10 Myr, t Time since SF began
“‘continuity bursty”’ N 6 bins (5 fitted parameters) First bin 0 – 10 Myr, SF begins at z = 20,
non-parametric SFH others distributed equally in log10 lookback time
d Student’s-t: Ratio of SFR in adjacent bins, coupled by
Charlot & Fall (2000) n clipped normal: Power-law slope of attenuation curve (A )
Dust Law (0.3, 2.5) For Calzetti et al. (2000) n
clipped normal: AV,<10Myr/AV ratio between young and old stars
(1, 3)
Salim et al. (2018) Dust Law clipped normal: Deviation from Calzetti et al. (2000) slope
(-0.3, 0.3)
uniform (0, 5) Strength of 2175Å bump
BPASS SPS Model No additional components
Uniform A Prior No additional components
Uniform Z⋆ Prior No additional components